Circular No. 6091

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

COMET MACHHOLZ (1994r) Donald E. Machholz, reports his visual discovery of a comet. The

following observations are available:

1994 UT  α_{2000}  δ_{2000}  m_{1} Observer
Oct. 8.505  8^{h}42.^{m}0  +55^{o}23′  11.5 Machholz
8.81957 8 41 44.^{s}72 +55 25 29′′.2 16.5 Tichá
8.82142 8 41 44.63 +55 25 29.6  _{”}
8.82392 8 41 44.46 +55 25 30.2  _{”}

D. E. Machholz (Colfax, CA). 0.25-m f/3.8 reflector, 36×. Comet diffuse

with weak condensation. J. Tichá and M. Tichý (Kleť Observatory). 0.57-m f/5.2 reflector + CCD

+ R filter. Measurer M. Tichý. Object slightly diffuse.

KS 1730−312

M. Gilfanov, E. Churazov, J. Ballet and E. Jourdain, on behalf of Granat team, report: “In the course of observations of the Galactic Center region on Sept. 26–27 Granat/SIGMA marginally detected KS 1730−312 (cf. IAUC 6083) with a 40–150-keV flux of 40 ± 15 mCrab. During subsequent observations on Sept. 28–30 the hard x-ray emission from the source was below SIGMA sensitivity (2σ limit 30 mCrab). The fast increase of the flux in the standard x-ray band observed by TTM (cf. IAUC 6088) was thus accompanied by a decrease of the hard x-ray flux, this behavior being consistent with dramatic steepening of the broad-band source spectrum observed by both TTM and SIGMA.”

PERIODIC COMET HARRINGTON (1994g) Total visual magnitude estimates: Aug. 7.39 UT, 13.1 (A. Hale, near

Chaparral, NM, 0.41-m reflector); 14.45, 13.2 (Hale); Sept. 7.03, 14.1 (K. Sárneczky, Ráktanya, Hungary, 0.44-m reflector); 9.42, 12.9 (Hale).

PERIODIC COMET BROOKS 2 (1994j) Total visual magnitude estimates by Hale: Aug. 14.46 UT, 13.8; Sept.

9.40, 13.9; 16.44, 13.8.

1994 October 8 Brian G. Marsden